Circular No. 8967

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 Q2 (ORY) An apparently asteroidal object discovered by Michel Ory (Delemont,

Switzerland, on CCD images obtained with a 0.61-m f/3.9 reflector at Vicques; discovery observation tabulated below), which was posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found by other CCD observers to be cometary. A. Knöfel (Schoenbrunn, Germany, 0.5-m reflector, Aug. 28.0 UT) reports an apparent faint coma. L. Buzzi (Varese, Italy, 0.60-m reflector, Aug. 28.1) notes a 12′′ coma elongated to the southwest, adding that his inspection of stacked astrographic images taken by R. Holmes (Charleston, IL, U.S.A., Aug. 28.28) reveals a coma ∼ 20′′ wide that is elongated in p.a. 240^{o}. A. C. Gilmore and P. M. Kilmartin remark that two stacked 120-s CCD frames taken with the 1.0-m f/7.7 reflector at Mt. John on Aug. 28.6 show a condensed, circular coma of diameter 14′′ with no tail.

2008 UT  α_{2000}  δ_{2000}  Mag.
Aug. 27.09470  1^{h}32^{m}40.^{s}74  +3^{o}04′17′′.2  17.6

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2008-Q51.

T = 2008 Oct. 23.018 TT  ω = 334.^{o}141
e = 0.58310  Ω = 59.478 }2000.0
q = 1.37046 AU  i = 2.814

a = 3.28723 AU n^{o} = 0.165370 P = 5.96 years

η CARINAE

A. Damineli — on behalf of a large team monitoring η Car spectroscopicaly at the Observatories of Pico dos Dias (Brazil), Casleo (Argentina), Las Campanas, and European Southern Observatory, La Silla — reports measurements in the He II 468.6-nm emission line. The equivalent width was 0.009 nm at epochs 2007.3 and 2007.6, 0.012 nm at 2008.3, and 0.021 nm at 2008.6 — repeating very well the past cycle, as reported by Steiner and Damineli (2004, Ap.J. 612, L133). The fact that the line-intensity curve follows the x-ray light curve indicates that the He II emitting region is associated with the wind-wind collision shock and not driven by a sudden shell ejection. For related information, see also Damineli et al. 2008, MNRAS 386, 2330.

2008 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green